OpenStudy (anonymous):

x − 10,000 =3/4 x What is the area of Cuba, in square miles? solve for x I do not want the answer just a little help on how to solve...

OpenStudy (zzr0ck3r):

Add 10000 to both sides then subtract 3/4X FROM BOTH SIDES AND TELL ME WHAT U GET

OpenStudy (zzr0ck3r):

Sorry forcaps

OpenStudy (zzr0ck3r):

Or subtract x from both sides

OpenStudy (anonymous):

x=3/4x + 10,000 x-3/4x=10,000 ?

OpenStudy (zzr0ck3r):

3/4x - x = -10000

OpenStudy (zzr0ck3r):

sorr I was on a tablet...

OpenStudy (zzr0ck3r):

so yes your right x-3/4x=10,000 then 1x - 3/4x = ?

OpenStudy (zzr0ck3r):

1/4x = 10000 x = 40000